OpenStudy (anonymous):

m + 3n = 10 m = n − 2

OpenStudy (anonymous):

Use substituion M value is given so we plug equation 2 into equation 1 \[\huge~\rm~n-2+3n=10\] from here solve for n

OpenStudy (anonymous):

N=4

OpenStudy (anonymous):

N=3

OpenStudy (anonymous):

n-2+3n = 10 add like terms. n+3n-2 = 10 Move -2 to the other side. 4n-2+2 = 10+2 4n = 12 4n/4 = 12/4 n = 3